# Privileged Information Distillation for Language Models

### v1

@@ 6 Out of Domain Experiments (OOD)
We find that for both Qwen3 models, $\pi$-Distill variants consistently outperform the base model and standard RL holding true for OPSD under more capable models. When compared to SFT w/ CoT + RL, we observe differences based on model size. On Qwen3-4B, SFT w/ CoT + RL is consistently the top performer with OPSD showing significant degradation. For Qwen3-8B, however, we find that $\alpha=0$ and $\alpha=0.5$ variants of $\pi$-Distill as well as OPSD can significantly outperform SFT w/ CoT + RL. For $\pi$-Distill, this aligns with our findings in section˜5. Where for Qwen3-4B not all $\pi$-Distill variants show substantial improvements over SFT w/ CoT + RL, whereas for Qwen3-8B, the opposite trend is observed. We observe a similar trend for OPSD, where it shows significant degradation on Qwen3-4B, while on Qwen3-8B it shows significant improvements. We attribute this to the possibility that on smaller models OPSD may overfit the teacher supervision for the task, while the more potent reasoners can provide more generalizable feedback. Overall, results on both $\pi$-Distill and OPSD imply that explicit supervision from frontier model’s CoT is more important for smaller models, where stronger reasoners benefit from being more on-policy via self-generated CoT. Additionally, we observe deterioration in R1-Distill-Llama-8B under SFT w/ CoT + RL, where performance drops below the base model. While $\pi$-Distill and OPSD do not improve over the base model in this specific case, they avoid significant deterioration. Finally, raw RL consistently exhibits degradation relative to the base model across all evaluations. Overall, these results indicate that $\pi$-Distill is highly effective for distillation in OOD settings when CoT traces are occluded, preventing performance from regressing below the base model. Moreover, both $\pi$-Distill and OPSD show significant improvements as model size scales.
Takeaways: OOD Generalization 1. Scaling Benefits for Stronger Reasoners: While SFT w/ CoT + RL is effective for Qwen3-4B, $\pi$-Distill and OPSD significantly outperform SFT w/ CoT + RL on Qwen3-8B. This suggests that as model capacity grows transfer can be enhanced by staying more on-policy. 2. Prevention of RL Degradation: Across all benchmarks, standard RL consistently performs worse than the base model. In contrast, OPSD avoids degradation in stronger reasoners, with $\pi$-Distill avoiding it in all cases.
@@ 7.1 What matters for $\pi$-Distill
Figure˜5 displays held-out performance on both $\tau$-Bench Retail and Travel Planner across the types of PI defined in § 4.2. Each subplot maps final performance against the initial divergence $D_{\mathrm{KL}}({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\text{base}}}\,\|\,{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\text{base}}})$, with accompanying bar plots showing the utility $\Delta$ for each type of PI. The most prominent pattern is that as the initial KL increases, final performance tends to decrease, though the full picture requires a more nuanced, $\alpha$-dependent analysis.
@@
Further, we find that PI utlity $\Delta$ can play a large role in the success of this variant. For instance, on Travel Planner with Qwen3-8B, $\alpha=0$ underperforms significantly because the PI provides negative utility ($\Delta<0$), actively degrading performance. Furthermore, as seen in figure˜6, we find that $\Delta_{\max}\approx 0$ in these failure cases, confirming that the algorithm cannot extract value from the PI when the signal itself offers no initial advantage over the base model. On the other hand, we find that when the teacher-student KL is low, student-only training effectively transfers knowledge back onto the teacher. This allows the teacher to learn how to leverage PI even though it is not being directly trained. We observe this behavior in Qwen3-8B on Travel Planner when using only tool calls as PI. While the initial $\Delta<0$, the low KL allows transfer to occur from the student to the teacher and enables the teacher to sample improved traces.
@@ 7.2 What Matters for OPSD
Figure 8: Performance and Stability Analysis of OPSD. We compare held-out performance on $\tau$-Bench (top row) and Travel Planner (bottom row) across three base models and three PI types (colors). The scatter plots map final scores against the student-teacher KL divergence ($D_{\text{KL}}({\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}}\|{\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}})$), while the bar-charts display the PI utility ($\Delta$). Key observations: (1) Unlike $\pi$-Distill, higher KL are not always detrimental, rather information richness of PI is most important, finding (Tool Calls & Arguments) often performs best (e.g., all results on Travel Planner and Qwen3-4B on $\tau$-Bench ). (2) Excessive KL can override positive utility (note Qwen3-8B on $\tau$-Bench, where $\Delta>0$ but the high KL degrades performance). (3) R1-Distill-Llama-8B consistently struggles, which we attribute to either extreme KL divergence ($\tau$-Bench) or negative PI utility (TravelPlanner).
@@
Interestingly, although we find that high-information PI is generally best suited for transfer, we observe exceptions. In particular, for Qwen3-8B on $\tau$-Bench, Tool Calls & Arguments performs the worst and exhibits the highest KL. In this case, we observe that $\Delta_{\max}$ is negative, indicating that the reverse-KL penalty can inhibit training. Additionally, regarding the failure on R1-Distill-Llama-8B, while this could be due to the SFT phase prior to OPSD, we find that on $\tau$-Bench the KL for both PI types is exceptionally high. While on Travel Planner, the PI utlity, $\Delta$ is low or negative, possibly explaining the diminished results. Overall, we find that the strongest predictor of performance for OPSD is the information content of the PI. While using the richest form of PI can lead to the best results, careful consideration is required to ensure the KL does not become excessive and hinder training.
@@ 9 Related Work
Recent work increasingly frames LM reasoning as a latent-variable inference problem (hu2024amortizingintractableinferencelarge; sordoni2023jointpromptoptimizationstacked; luo2025languagemodelslearnverbal; li2025startselftaughtreasonertools). Within this framework, the most popular approach is STaR (li2025startselftaughtreasonertools), which uses a hint when the model is unable to correctly answer the question and then uses SFT to fit onto the generated reasoning trace. This is similar to student-only training in $\pi$-Distill ($\alpha=0$), where rather than a KL-regularized off-policy RL objective, SFT is used and the teacher is not trained. Most similar to $\pi$-Distill is the work by (zhou2025variationalreasoninglanguagemodels), which proposes variational reasoners. This approach can be seen as a modified version of STaR, where both teacher and student are being trained iteratively. The main difference between variational reasoning and our work is they assume access to oracle answers and perform an iterative version of variational EM using separate parameters for the teacher and student. In comparison, we simplify the training objective by allowing the teacher and student to share parameters and do not assume access to ground truth answers. We compare against a similar setup with minor modifications due to not having access to ground truth solutions in App. A.2.
@@
Complementary work focuses on using privileged or auxiliary signals to enable learning in hard regimes. (chen2025nudgingboundariesllmreasoning) injects self-generated high-level hints into online RL to overcome zero-reward exploration barriers, while (qu2026popelearningreasonhard) uses privileged oracle solutions as structured on-policy exploration signals for hard reasoning tasks. Both these algorithms can be viewed as training the teacher to use PI with implicit transfer via parameter sharing. Additionally, other lines of work explore distilling certain skills/behaviors in the model via contextualized sampling (didolkar2025metacognitivereuseturningrecurring; didolkar2024metacognitivecapabilitiesllmsexploration). We believe our proposed methods could be used to distill desired skills/behaviors into model weights in a more effective manner compared to traditional SFT.
@@ 10 Limitations & Future Work
While we show that both $\pi$-Distill and OPSD are effective algorithms when CoT is not available, all our PI is transformed from frontier model traces. Exploring how these algorithms perform without access to a frontier model or ground truth answers is an interesting direction. Additionally, our experiments are limited to models $\leq 8$B parameters. Further scaling of both $\pi$-Distill and OPSD can shine light on additional important factors to enable these algorithms. Moreover, our analysis in Section˜7 is limited to observational studies, where we do not systematically control for all variables, rather observe properties of different configurations. A study that systematically controls for these factors could help in mitigating failure cases and obtaining the most useful PI.
@@ 12 Acknowledgments
We thank and credit Siddarth Venkatraman for the idea to view learning from PI from a variational perspective which led to the final version of $\pi$-Distill. We thank Vedant Shah with helpful discussion and proof reading initial drafts of the paper. Additionally we thank, Michael Noukhovitch for valuable feedback that substantially aided the final draft of this work. We also thank Alexandre Piché for their valuable discussions.
@@ 13 Author Contributions
Emiliano Penaloza proposed the idea of leveraging PI, led the project, implemented all algorithms and the RL environments, conducted most experiments and was the primary writer of the paper.
@@
Alexandre Lacoste advised the project providing valuable insights to the goal and purpose of leveraging PI, helped improve experimental rigor and analysis, and assisted in paper writing.
Laurent Charlin advised the project helping influence the framing and perspective of the project, provided help with writing with a substantial portion of the introduction written by them.
@@ Appendix A Connections To Variational EM
In this Appendix, we derive $J_{\pi-\text{Distill}}$ using a variational perspective. Specifically we show how $J_{\pi-\text{Distill}}$ can be seen as a joint EM algorithm where E and M are done simultaneously.
Traditional variational-EM aims to fit a parameterized distribution $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} onto a target posterior $\pi^{*}$. An assumption this framework makes is that using $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} to approximate $\pi^{*}$ may be suboptimal, rather leveraging an approximate posterior $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})$} conditioned on additional information can make the approximation easier. This is true when $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}$} can sample from high-reward modes that $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} cannot, therefore making optimization via RL impossible for $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}. We visualize the procedure in Figure˜10
@@ A.1 Variational EM
Traditionally, one would either first fit $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}$} to convergence and use it to fit $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}$}, or alternate between training the teacher and student. Regardless, this process requires two separate models and can be difficult to optimize. Our implementation is a modified version of zhou2025variationalreasoninglanguagemodels, where the modifications are specific to our multi-turn setting where no ground truth is provided.
@@
Figure 11: Performance of Sequential EM variants on $\tau$-Bench Retail (Qwen3-8B). We compare these methods against SFT baselines. While standard EM RFT (light blue) fails to match SFT w/ CoT (yellow), replacing RFT with off-policy RL (EM Off-Pol, purple) successfully closes allows for EM to outperform SFT w CoT. The strongest performance comes from $J_{\text{Teacher}}(\theta)$ (dark blue), demonstrating that parameter sharing between the teacher and student yields the most effective transfer. Shaded regions indicate standard error across 3 seeds.
Our initial efforts focused on making variational EM work, similar in spirit to zhou2025variationalreasoninglanguagemodels. We first document initial ideas that were unsuccessful, then we discuss the best version of EM we could find being a sequential implementation. Finally we discuss how these experiments ultimately motivated the design decisions for $\pi$-Distill.
@@
Our following efforts focused on training the teacher online using LABEL:eq:j_teacher. We try this with alternating loops between student and teacher training, but find this to be ineffective. As we found it hard to avoid the student and teacher collapsing onto each other using this setup. Ultimately, we found a sequential version of EM worked best, though it still requires specific adjustments to achieve optimal performance.
@@
We turn to Sequential EM after ruling out offline and alternating approaches. However, we find standard implementations remain unreliable for complex agentic tasks. Here we analyze the limitations of standard EM and how they ultimately motivated $\pi$-Distill.
For this analysis we use $\tau$-Bench Retail with Qwen3-8B using self-generated hints to train the teacher as this configuration performs best (see § 7). We compare sequential EM, where we first train the teacher and then perform SFT on successful trajectories (Rejection Fine-tuning or RFT), against SFT with and without CoT. We note that here it would be sufficient to outperform SFT without CoT as we assume no access to internal CoT of the frontier model, thus in practice sequential EM could replace the traditional SFT phase before RL.
We report results in Figure˜11. First, we observe that standard Sequential EM outperforms naive SFT but lags behind training with CoT. We attribute the performance gap to the lack of negative feedback on failed trajectories. It only mimics the teacher. To address this, we replaced the RFT step with clipped off-policy RL. We find this change enables sequential EM to outperform SFT with CoT.

### v3

@@ 6 Out of Domain Experiments (OOD)
We find that for both Qwen3 models, $\pi$-Distill variants consistently outperform the base model and standard RL holding true for OPSD under more capable models. When compared to SFT w/ CoT + RL, we observe differences based on model size. On Qwen3-4B, SFT w/ CoT + RL is consistently the top performer with OPSD showing significant degradation. For Qwen3-8B, however, we find that $\alpha=0$ and $\alpha=0.5$ variants of $\pi$-Distill as well as OPSD can significantly outperform SFT w/ CoT + RL. For $\pi$-Distill, this aligns with our findings in section˜5. Where for Qwen3-4B not all $\pi$-Distill variants show substantial improvements over SFT w/ CoT + RL, whereas for Qwen3-8B, the opposite trend is observed. We observe a similar trend for OPSD, where it shows significant degradation on Qwen3-4B, while on Qwen3-8B it shows significant improvements. We attribute this to the possibility that on smaller models, OPSD may overfit the teacher supervision for the task, while the more potent reasoners can provide more generalizable feedback. Overall, results on both $\pi$-Distill and OPSD imply that explicit supervision from frontier model’s CoT is more important for smaller models, where stronger reasoners benefit from being more on-policy via self-generated CoT. Additionally, we observe deterioration in R1-Distill-Llama-8B under SFT w/ CoT + RL, where performance drops below the base model. While $\pi$-Distill and OPSD do not improve over the base model in this specific case, they avoid significant deterioration. Finally, raw RL consistently exhibits degradation relative to the base model across all evaluations. Overall, these results indicate that $\pi$-Distill is highly effective for distillation in OOD settings when CoT traces are occluded, preventing performance from regressing below the base model. Moreover, both $\pi$-Distill and OPSD show significant improvements as model size scales.
Takeaways: OOD Generalization 1. Scaling Benefits for Stronger Reasoners: While SFT w/ CoT + RL is effective for Qwen3-4B, $\pi$-Distill and OPSD significantly outperform SFT w/ CoT + RL on Qwen3-8B. This suggests that as model capacity grows, transfer can be enhanced by staying more on-policy. 2. Prevention of RL Degradation: Across all benchmarks, standard RL consistently performs worse than the base model. In contrast, OPSD avoids degradation in stronger reasoners, with $\pi$-Distill avoiding it in all cases.
@@ 7.1 What matters for $\pi$-Distill
Figure˜5 displays held-out performance on both $\tau$-Bench Retail and Travel Planner across the types of PI defined in Section˜4.2. Each subplot maps final performance against the initial divergence $D_{\mathrm{KL}}({\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\text{base}}}\,\|\,{\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\text{base}}})$, with accompanying bar plots showing the utility $\Delta$ for each type of PI. The most prominent pattern is that as the initial KL increases, final performance tends to decrease, though the full picture requires a more nuanced, $\alpha$-dependent analysis.
@@
Further, we find that PI utility $\Delta$ can play a large role in the success of this variant. For instance, on Travel Planner with Qwen3-8B, $\alpha=0$ underperforms significantly because the PI provides negative utility ($\Delta<0$), actively degrading performance. Furthermore, as seen in figure˜6, we find that $\Delta_{\max}\approx 0$ in these failure cases, confirming that the algorithm cannot extract value from the PI when the signal itself offers no initial advantage over the base model. On the other hand, we find that when the teacher-student KL is low, student-only training effectively transfers knowledge back onto the teacher. This allows the teacher to learn how to leverage PI even though it is not being directly trained. We observe this behavior in Qwen3-8B on Travel Planner when using only tool calls as PI. While the initial $\Delta<0$, the low KL allows transfer to occur from the student to the teacher and enables the teacher to sample improved traces.
@@ 7.2 What Matters for OPSD
Figure 8: Performance and Stability Analysis of OPSD. We compare held-out performance on $\tau$-Bench (top row) and Travel Planner (bottom row) across three base models and three PI types (colors). The scatter plots map final scores against the student-teacher KL divergence ($D_{\text{KL}}({\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}}\|{\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}})$), while the bar-charts display the PI utility ($\Delta$). Key observations: (1) Unlike $\pi$-Distill, higher KL are not always detrimental, rather, information richness of PI is most important, finding (Tool Calls & Arguments) often performs best (e.g., all results on Travel Planner and Qwen3-4B on $\tau$-Bench ). (2) Excessive KL can override positive utility (note Qwen3-8B on $\tau$-Bench, where $\Delta>0$ but the high KL degrades performance). (3) R1-Distill-Llama-8B consistently struggles, which we attribute to either extreme KL divergence ($\tau$-Bench) or negative PI utility (TravelPlanner).
@@
Interestingly, although we find that high-information PI is generally best suited for transfer, we observe exceptions. In particular, for Qwen3-8B on $\tau$-Bench, Tool Calls & Arguments performs the worst and exhibits the highest KL. In this case, we observe that $\Delta_{\max}$ is negative, indicating that the reverse-KL penalty can inhibit training. Additionally, regarding the failure on R1-Distill-Llama-8B, while this could be due to the SFT phase prior to OPSD, we find that on $\tau$-Bench the KL for both PI types is exceptionally high. While on Travel Planner, the PI utility, $\Delta$ is low or negative, possibly explaining the diminished results. Overall, we find that the strongest predictor of performance for OPSD is the information content of the PI. While using the richest form of PI can lead to the best results, careful consideration is required to ensure the KL does not become excessive and hinder training.
@@ 9 Related Work
Recent work increasingly frames LM reasoning as a latent-variable inference problem (hu2024amortizingintractableinferencelarge; sordoni2023jointpromptoptimizationstacked; luo2025languagemodelslearnverbal; li2025startselftaughtreasonertools). Within this framework, the most popular approach is STaR (li2025startselftaughtreasonertools), which uses a hint when the model is unable to correctly answer the question and then uses SFT to fit onto the generated reasoning trace. This is similar to student-only training in $\pi$-Distill ($\alpha=0$), where, rather than a KL-regularized off-policy RL objective, SFT is used, and the teacher is not trained. Most similar to $\pi$-Distill is the work by (zhou2025variationalreasoninglanguagemodels), which proposes variational reasoners. This approach can be seen as a modified version of STaR, where both teacher and student are being trained iteratively. The main difference between variational reasoning and our work is that they assume access to oracle answers and perform an iterative version of variational EM using separate parameters for the teacher and student. In comparison, we simplify the training objective by allowing the teacher and student to share parameters and do not assume access to ground truth answers. We compare against a similar setup with minor modifications due to not having access to ground truth solutions in App. A.2.
@@
Complementary work focuses on using privileged or auxiliary signals to enable learning in hard regimes. chen2025nudgingboundariesllmreasoning injects self-generated high-level hints into online RL to overcome zero-reward exploration barriers, while qu2026popelearningreasonhard uses privileged oracle solutions as structured on-policy exploration signals for hard reasoning tasks. Both these algorithms can be viewed as training the teacher to use PI with implicit transfer via parameter sharing. Additionally, other lines of work explore distilling certain skills/behaviors in the model via contextualized sampling (didolkar2025metacognitivereuseturningrecurring; didolkar2024metacognitivecapabilitiesllmsexploration; yang-etal-2025-distilling; qu2025rladtrainingllmsdiscover). We believe our proposed methods could be used to distill desired skills/behaviors into model weights in a more effective manner compared to traditional SFT.
Context Distillation.
Both our proposed methods can be seen as a form of context distillation. This growing line of work studies how to train models to internalize context-dependent reasoning so that the benefits of rich contexts (e.g., instructions, demonstrations, auxiliary computations) are distilled into the model’s weights (caccia2025trainingplugnplayknowledgemodules; snell2022learningdistillingcontext; huang2022incontextlearningdistillationtransferring). Other works analyze this problem by performing SFT answers they give when having access to context snell2022learningdistillingcontext. Moreover, other works also try to encapsulate such knowledge into adapters that can be plugged and played at test time (caccia2025trainingplugnplayknowledgemodules).
@@ 10 Limitations & Future Work
While we show that both $\pi$-Distill and OPSD are effective algorithms when CoT is not available, all our PI is transformed from frontier model traces. Exploring how to efficiently obtain useful PI when neither frontier model actions nor ground-truth answers are available is an interesting direction. Concurrent work by hubotter2026reinforcementlearningselfdistillation explores this direction by mining PI through self-reflection. Additionally, our experiments are limited to models with $\leq 8$B parameters. Further scaling of both $\pi$-Distill and OPSD can shine light on additional important factors to enable these algorithms. Moreover, our analysis in section˜7 is limited to observational studies, where we do not systematically control for all variables, but rather observe properties of different configurations. A study that systematically controls for these factors could help in mitigating failure cases and obtaining the most useful PI.
@@ 12 Acknowledgments
We thank and credit Siddarth Venkatraman for the idea to view learning from PI from a variational perspective, which led to the final version of $\pi$-Distill. We thank Vedant Shah for helpful discussion and proof reading initial drafts of the paper. Additionally, we thank Michael Noukhovitch for valuable feedback that substantially aided the final draft of this work. We also thank Alexandre Piché for their valuable discussions.
@@ 13 Author Contributions
Emiliano Penaloza proposed the idea of leveraging PI, led the project, implemented all algorithms and the RL environments, conducted most experiments, and was the primary writer of the paper.
@@
Alexandre Lacoste advised the project, providing valuable insights to the goal and purpose of leveraging PI, helped improve experimental rigor and analysis, and assisted in paper writing.
Laurent Charlin advised the project, helping influence the framing and perspective of the project, and provided help with writing, with a substantial portion of the introduction written by them.
@@ Appendix A Connections To Variational EM
In this Appendix, we derive $J_{\pi-\text{Distill}}$ using a variational perspective. Specifically, we show how $J_{\pi-\text{Distill}}$ can be seen as a joint EM algorithm where E and M are done simultaneously.
Traditional variational-EM aims to fit a parameterized distribution $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} onto a target posterior $\pi^{*}$. An assumption this framework makes is that using $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} to approximate $\pi^{*}$ may be suboptimal, rather, leveraging an approximate posterior $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}(\mathbf{o}\mid s,\mathbf{I})$} conditioned on additional information can make the approximation easier. This is true when $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}$} can sample from high-reward modes that $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$} cannot, therefore making optimization via RL impossible for $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}$}. We visualize the procedure in Figure˜10
@@ A.1 Variational EM
Traditionally, one would either first fit $\color[rgb]{0.390625,0.2578125,0.828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.390625,0.2578125,0.828125}\pi^{\mathrm{T}}_{\theta}$} to convergence and use it to fit $\color[rgb]{0.90234375,0.59765625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.90234375,0.59765625,0}\pi^{\text{S}}_{\theta}$}, or alternate between training the teacher and student. Regardless, this process requires two separate models and can be difficult to optimize. Our implementation is a modified version of zhou2025variationalreasoninglanguagemodels, where the modifications are specific to our multi-turn setting, where no ground truth is provided.
@@
Figure 11: Performance of Sequential EM variants on $\tau$-Bench Retail (Qwen3-8B). We compare these methods against SFT baselines. While standard EM RFT (light blue) fails to match SFT w/ CoT (yellow), replacing RFT with off-policy RL (EM Off-Pol, purple) successfully allows for EM to outperform SFT w CoT. The strongest performance comes from $J_{\text{Teacher}}(\theta)$ (dark blue), demonstrating that parameter sharing between the teacher and student yields the most effective transfer. Shaded regions indicate standard error across 3 seeds.
Our initial efforts focused on making variational EM work, similar in spirit to zhou2025variationalreasoninglanguagemodels. We first document initial ideas that were unsuccessful, then we discuss the best version of EM we could find, being a sequential implementation. Finally, we discuss how these experiments ultimately motivated the design decisions for $\pi$-Distill.
@@
Our following efforts focused on training the teacher online using Section˜A.1. We try this with alternating loops between student and teacher training, but find this to be ineffective. As we found it hard to avoid the student and teacher collapsing onto each other using this setup. Ultimately, we found a sequential version of EM worked best, though it still requires specific adjustments to achieve optimal performance.
@@
We turn to Sequential EM after ruling out offline and alternating approaches. However, we find standard implementations remain unreliable for complex agentic tasks. Here, we analyze the limitations of standard EM and how they ultimately motivated $\pi$-Distill.
For this analysis, we use $\tau$-Bench Retail with Qwen3-8B using self-generated hints to train the teacher as this configuration performs best (see § 7). We compare sequential EM, where we first train the teacher and then perform SFT on successful trajectories (Rejection Fine-tuning or RFT), against SFT with and without CoT. We note that here it would be sufficient to outperform SFT without CoT as we assume no access to the internal CoT of the frontier model, thus in practice, sequential EM could replace the traditional SFT phase before RL.
We report results in Figure˜11. First, we observe that standard Sequential EM outperforms naive SFT but lags behind training with CoT. We attribute the performance gap to the lack of negative feedback on failed trajectories. It only mimics the teacher. To address this, we replaced the RFT step with clipped off-policy RL. We find that this change enables sequential EM to outperform SFT with CoT.